Caspian Sea Drinks is considering the production of a diet drink. The expansion of the plant and the purchase of the equipment necessary to produce the diet drink will cost $27.00 million. The plant and equipment will be depreciated over 10 years to a book value of $1.00 million, and sold for that amount in year 10. Net working capital will increase by $1.07 million at the beginning of the project and will be recovered at the end. The new diet drink will produce revenues of $9.36 million per year and cost $2.23 million per year over the 10-year life of the project. Marketing estimates 11.00% of the buyers of the diet drink will be people who will switch from the regular drink. The marginal tax rate is 25.00%. The WACC is 14.00%. Find the IRR (internal rate of return).
Answer #1

calculation of depreciation

Depreciation = [Cost - Book value at the end] / useful life = (27 - 1) / 10 = $2.6 million

total outflow = purchase value of equipment + Net Working Capital

= 27 + 1.07 = $28.07 million

calculating operating cash flows:

Revenue from diet drink = 9,360,000

less : costs = 2,230,000

less: Depreciation = 2,600,000

profit before tax = 4,530,000

less : tax @ 25% = 1,132,500

profit after tax = 3,397,500

Add : Depreciation = 2,600,000

Operating Cash Flow = 5,997,500

Operating cash flow will be same every year except for last year because in last year there will be recovery of working capital

so cash flow from years 1 to 9 = $5,997,500

in year 10 it is = 5,997,500 + 1070000 = $7,067,500

IRR is the rate where NPV of the project is 0

year Net cash flow
0 -28,070,000
1 5997500
2 5997500
3 5997500
4 5997500
5 5997500
6 5997500
7 5997500
8 5997500
9 5997500
10 7067500
IRR 17.09%

Using Excel ( formula is =IRR(select all above cells) i.e., from year 0 to 10) ) IRR comes to 17.09%

NOTE: The percentage of buyers shifting from regular to diet drink is irrelevant for this project.